OpenStudy (anonymous):

The length of a rectangle is 3yrds longer than its width. If the perimeter of the rectangle is 46yd , find its area.

OpenStudy (anonymous):

Perimeter formula P= 2L +2W length is 3 + longer L = 3+W W = W P = 46 with this information you can begin solving all you just need to do is plug in the information you should get and answer followed then you can use the area formula after finding length and width Area formula A=LW

OpenStudy (anonymous):

do you understand how to do these types of problems BTW @ashleynicolee